# SCoTER: Structured Chain-of-Thought Transfer for Enhanced Recommendation

### v4

@@ Appendix B Manual CoT Templates
Prompt: Inferring Recommendation Paths from User Behavior You are an expert in recommendation algorithms. Based on the user’s historical behavior data, please infer a reasonable analysis and recommendation path, and predict 20 different product features that the user may be interested in the future. Requirements: (1) First output <cot_path>, with a reasoning chain of at least two steps, each step $\leq$8 words, connected with ”-¿”. (2) <reason> must be placed after <cot_path>, and the content must strictly follow each step in <cot_path> to expand the reasoning one by one. (3) Finally output 20 recommended product feature descriptions. Output format requirements: <cot_path> [Summarize the core reasoning path here: Step 1 -¿ Step 2-¿ $\cdots$ -¿ Item recommendation] </cot_path> <reason> [Explain the reasoning logic step by step according to the steps in cot_path] </reason> <recommendations> <item>Detailed feature description of product 1, including category, brand, function, characteristics, etc.</item> <item>Detailed feature description of product 2, including category, brand, function, characteristics, etc.</item> $\cdots$ <item>Detailed feature description of product 20, including category, brand, function, characteristics, etc.</item> </recommendations>

### v5

@@ Appendix F CoT Templates
Prompt: Inferring Recommendation Paths from User Behavior You are an expert in recommendation algorithms. Based on the user’s historical behavior data, please infer a reasonable analysis and recommendation path, and predict 20 different product features that the user may be interested in the future. Requirements: (1) First output <cot_path>, with a reasoning chain of at least two steps, each step $\leq$8 words, connected with ”-¿”. (2) <reason> must be placed after <cot_path>, and the content must strictly follow each step in <cot_path> to expand the reasoning one by one. (3) Finally output 20 recommended product feature descriptions. Output format requirements: <cot_path> [Summarize the core reasoning path here: Step 1 -¿ Step 2-¿ $\cdots$ -¿ Item recommendation] </cot_path> <reason> [Explain the reasoning logic step by step according to the steps in cot_path] </reason> <recommendations> <item>Detailed feature description of product 1, including category, brand, function, characteristics, etc.</item> $\cdots$ <item>Detailed feature description of product 20, including category, brand, function, characteristics, etc.</item> </recommendations>
@@
Mined Pattern (GVM-Discovered)
Unlike manual templates based on general heuristics, GVM discovers patterns that capture domain-specific logic (e.g., separating stable vs. recent interests in Beauty, or identifying complementary gear in Sports).
Beauty (Universal Style):
<cot_path >Analyze sequential purchase history -¿ Distinguish stable vs. recent interests -¿ Expand via shared item attributes -¿ Predict high-confidence features -¿ Item Recommendation </cot_path>
Instruments:
<cot_path >Analyze the purchase pattern -¿ Music interest inference -¿ Instrument focus identification -¿ Accessory expansion -¿ Feature prediction </cot_path>
Sports:
<cot_path >Purchase pattern recognition -¿ Outdoor/Athletic interest mapping -¿ Complementary gear expansion-¿ Item Recommendation </cot_path>
Yelp:
<cot_path >Purchase analysis -¿ Cuisine/Location interest mapping -¿ Category expansion -¿ Feature prediction </cot_path>
Prompt: Mining Optimal Patterns from CoT Clusters You are an expert in recommendation system reasoning patterns. Analyze the following CoT cluster and propose better reasoning paths. Cluster ID: x Performance: Rank 1 in K clusters $|$ {R@20 value} Recall@20 Representative CoT Paths: (1) <cot_path > $\cdots$ </cot_path> $\cdots$ (2) <cot_path > $\cdots$ </cot_path>   Analysis & Output: ### 1. Pattern Summary Core Logic: [One sentence describing the main reasoning approach] Key Steps: [The most important characteristics of this pattern] ### 2. Performance Assessment Why it works: [What makes this pattern achieve its current hit rate] Limitations: [Main weaknesses preventing higher performance] ### 3. Optimized Paths Best Path (each step $\leq$ 8 words): [Step1 -¿ Step2 -¿ … -¿ Stepk -¿ Item Recommendation] Improvement Logic: [How these address the identified limitations and why they should perform better] Requirements: Keep the CoT path clear and not overly complex
Figure 10. The prompt structure used in the Mine phase. The LLM acts as an expert analyst to abstract a single, optimal reasoning template from a cluster of high-performing candidate chains.
Example [Instruction Template]
You are an expert recommendation system. Your task is to analyze user purchase patterns and predict their next likely purchase based on their sequential purchase history.
Reasoning Pattern (High-level Steps): Analyze sequential purchase history $\to$ Distinguish stable vs. recent interests $\to$ Expand via shared item attributes $\to$ Predict high-confidence features $\to$ Item Recommendation Required Output Format: Start with Overall Analysis. Then provide detailed analysis for each step in the Reasoning Pattern. End with Summary. [Input Data]
Analyze the following user purchase history and provide a comprehensive
structured rationale following the exact format:
Purchase History in Chronological Order: ## Item 1: OPI Nail Lacquer, Simmer and Shimmer. Desc: OPI Burlesque Colors. ## Item 2: OPI Red Shatter Crackle Nail Polish. Desc: Red Shatter Nail Polish. ## Item 3: SKIN79 Prestige Beblesh Balm $B$ Cream. Desc: Whitening, wrinkles improvement… Diamond Powder… ## Item 4: WAWO 15 Color Professional Camouflage Concealer. Desc: Camouflage for blemishes, scars… perfect for under-eye area…   [Model Output: Structured Rationale] Overall Analysis: The user demonstrates a clear focus on detailed personal grooming. Their behavior suggests a preference for ”professional-grade” home beauty solutions, transitioning from expressive decoration to functional care. [Analyze sequential purchase history]: The sequence begins with vibrant hand aesthetics (Items 1-2), prioritizing visual impact. It then shifts chronologically towards corrective facial skincare (Items 3-4), indicating a move from purely decorative cosmetics to functional, skin-improving formulations. [Distinguish stable vs. recent interests]: Stable: A consistent demand for ”Salon-quality/Professional tools” (e.g., OPI brand, 15-color palette). Recent: The immediate focus has pivoted to ”Skin Health” and ”Restoration” (whitening, moisturizing), showing a specific intent to improve skin condition rather than just painting over it. [Expand via shared item attributes]: We synthesize the Location from the stable history (Hands) with the Function from the recent history (Therapeutic/Skin Softening). The expansion leads to ”Restorative Hand Treatments.” [Predict high-confidence features]: The next item should likely offer: ”Thermal/Heating” (for deep treatment), ”Moisturizing” (aligning with skincare logic), and ”Spa-like efficacy.” [Item Recommendation]: A professional-grade thermal paraffin wax bath or home spa kit. Rationale: This product category perfectly synthesizes the user’s profile. It addresses the ”Hand Care” interest established by the OPI purchases but elevates it to the ”Treatment/Care” level seen in the recent skincare purchases.
Figure 11. A complete example of data synthesis. The upper part shows the pattern-injected instruction and user history provided to the Teacher LLM. The lower part shows the structured rationale generated by the LLM.